# Are LLM Evaluators Really Narcissists? Sanity Checking Self-Preference Evaluations

### v1

@@ Abstract
Recent research has shown that large language models (LLM) favor own outputs when acting as judges, undermining the integrity of automated post-training and evaluation workflows. However, it is difficult to disentangle which evaluation biases are explained by narcissism versus general experimental confounds, distorting measurements of self-preference bias. We discover a core methodological confound which could reduce measurement error by 89.6%. Specifically, LLM evaluators may deliver self-preferring verdicts when the judge responds to queries which they completed incorrectly themselves; this would be true regardless of whether one of their responses is their own. To decouple self-preference signals from noisy outputs on hard problems, we introduce an Evaluator Quality Baseline, which compares the probability that a judge incorrectly votes for itself against the probability that it votes for an incorrect response from another model. Evaluating this simple baseline on 37,448 queries, only 51% of initial findings retain statistical significance. Finally, we turn towards characterizing the entropy of “easy” versus “hard” evaluation votes from LLM judges. Our corrective baseline enables future research on self-preference by eliminating noisy data from potential solutions. More widely, this work contributes to the growing body of work on cataloging and isolating judge-bias effects.
@@ 1 Introduction
As language models see increasing deployment as evaluators, raters, and proxies for human preference, researchers have observed consistent patterns of unreliability Gu et al. (2025); Li et al. (2024). These patterns are frequently categorized as biases Dietz et al. (2025); Ye et al. (2024) and are often interpreted through anthropomorphic tropes or linked to emerging theory-of-mind frameworks.
One notable example comes with the recent fascination with self-preference bias, or narcissism, of LLM evaluators Panickssery et al. (2024); Liu et al. (2024b). A suite of recent experimental findings claim that when language models are used to compare responses to a query, the distribution of votes will skew towards outputs generated by that same model. If true, this behavior would pose critical implications for safety and alignment research; this form of situational awareness Laine et al. (2023) risks contaminating post-training, guardrail construction, and monitoring efforts that bootstrap human preferences using LLM evaluators.
However, self-preference signals remain confounded by plausible baselines. For instance, early tests did not compare self-preference rates against ground truth accuracy, potentially mistaking a model’s accurate assessment of its own superior performance for illegitimate bias. Recent efforts now control for this by using oracle labels Chen et al. (2025a); Wataoka et al. (2025), but still fail to disentangle the relationship between task accuracy and evaluation accuracy; that is, the ability of a judge model to assign accurate preference ratings regardless of its own skill in a given domain. The capabilities a model exhibits on a given task may inform its preferences on that same task; specifically, one might imagine that an LLM-judge’s distribution of preferences will differ depending on if it could succeed on the query in question. If evaluations on these “harder” queries show greater uncertainty, then a model might choose its own, incorrect response at random. The inability to distinguish between a model’s general evaluative proficiency and an active self-bias undermines the causality of previously established results.
@@ 2.1 Self-Preference and Other Biases in LLM Evaluators
Literature on LLM evaluation has rapidly expanded in recent years, with numerous surveys cataloging the diverse methodologies and challenges (Zheng et al., 2023; Shen et al., 2023; Chen et al., 2024; Li et al., 2024, 2025a; Gu et al., 2025). A significant subset of this work focuses on the reliability and biases of LLM evaluators, identifying various failure modes including inconsistency, susceptibility to manipulation, and context sensitivity Koo et al. (2024); Ye et al. (2024); Dietz et al. (2025).
@@ 3.3 The Need for an Evaluator Quality Baseline
If the ILSP term controls bias, how can we characterize the distribution that produces it? All examples which fall into this subset are examples where a judge model produced a verifiably incorrect or relatively inferior response. This subset introduces a covariate on evaluations based on the judge’s task-related capability. We assert that this deficit in capability influences the voting behavior of the judge. This is analogous to a student’s performance on a multiple choice exam (pairwise preference), which may be lower-bounded by their performance in a free response setting (generative task). This student’s “multiple-choice” behavior may manifest as pure aleatory uncertainty or deference to shallow heuristics. Research on uncertainty quantification has shown that while evaluator errors do depend on task accuracy Kalai et al. (2025); Du et al. (2024); Sychev et al. (2025), the mechanisms which drive such errors vary by model, task, and context.
@@ 4.2 Reproducing Previous Experiments
4.2.1 Verifiable Outputs Chen et al. (2025a)
We reproduce the verifiable reasoning experiments of Chen et al. (2025a), which isolate self-preference on tasks with objectively correct answers. Datasets include MATH-500Hendrycks et al. (2021b), MBPP-Plus Austin et al. (2021), and MMLU Hendrycks et al. (2021a). Ground truth is established via execution-based evaluation for code (pass@1) and exact string matching for mathematics and multiple-choice responses. This setup enables clean separation of LSP from ILSP examples.
@@
4.2.2 Debiased Gold Judges Chen et al. (2025b)
We reproduce the DBG-scoring methodology of Chen et al. (2025b), which evaluates self-preference on open-ended tasks including helpfulness with AlpacaEval Li et al. (2023), truthfulness with TruthfulQA Lin et al. (2022), and translation quality with the German-English WMT19 dataset Barrault et al. (2019).
@@
4.2.3 Human Annotations Panickssery et al. (2024)
We reproduce the summarization experiments of Panickssery et al. (2024), which established a correlation between self-recognition and self-preference on news summarization datasets, namely CNN/DailyMail See et al. (2017) and XSUM Narayan et al. (2018). Similar to DBG-Score, ground-truth summaries are generated at-scale with a “debiased gold judge”—in this case, GPT 5.2—with a prompt first tuned by small-scale human validation before tested on a larger, holdout set. We take log-probabilities through the OpenAI API, following the hyperparameters used initially. Our proxy selection matches the outcome-based approach used for other experimental setups, except that we only take a subset of potential proxies due to costs of accruing gold-judge labels from proprietary models.
@@ 4.2.4 Authorship Obfuscation
Lastly, we replicate the setup of Mahbub and Feng (2025), which investigates self-preference under stylistic perturbation using answer extraction as the verifiability signal. This approach tests whether synonym replacement and paraphrasing eliminate self-preference by disrupting surface-level cues. This work tests self-preference on the multiple-choice reading comprehension dataset QualiTY Pang et al. (2022). Since reading comprehension responses are in multiple-choice format, we can gather ground truth labels by parsing the corresponding answer and outcome-match queries accordingly.
@@ 5.1 The Evaluator Quality Baseline Substantially Reduces Measured Self-Preference
Our baseline complements rather than contradicts prior findings: self-preference exists, but the prioritization of at-risk models has been misplaced by measurement error. Rankwise model bias orderings shift considerably after applying the Evaluator Quality Baseline. Using Alpaca Eval as an example, Llama 3.1-8B showed the highest initial bias rate of 49% before the baseline, and now shows just 3.2%—the second lowest. On the other hand, for code generation, Qwen2.5-72B went from the third-to-least to the highest rankwise bias relative to its baseline. For data-based mitigation strategies like steering vectors Roytburg et al. (2025); Ackerman and Panickssery (2024), having controlled measurements determines choice of mechanism design and training targets, enabling more effective interventions.
@@ Appendix A Experimental Configuration Tables
Since we are using a pairwise setting, to account for positional bias Pezeshkpour and Hruschka [2023] we prompt the judge twice, once where the judges prompt is first and once where the judges prompt appears second. We collect from both prompts logprobs and average them to get the probabilities for self.

### v2

@@ Abstract
Recent research has shown that large language models (LLMs) favor their own outputs when acting as judges, undermining the integrity of automated post-training and evaluation workflows. However, it is difficult to disentangle which evaluation biases are explained by narcissism versus general experimental confounds, distorting measurements of self-preference bias. We discover a core methodological confound which could reduce measurement error by 89.6%. Specifically, LLM evaluators may deliver self-preferring verdicts when the judge responds to queries which they completed incorrectly themselves; this would be true regardless of whether one of their responses is their own. To decouple self-preference signals from noisy outputs on hard problems, we introduce an Evaluator Quality Baseline, which compares the probability that a judge incorrectly votes for itself against the probability that it votes for an incorrect response from another model. Evaluating this simple baseline on 37,448 queries, only 51% of initial findings retain statistical significance. Finally, we turn towards characterizing the entropy of “easy” versus “hard” evaluation votes from LLM judges. Our corrective baseline enables future research on self-preference by eliminating noisy data from potential solutions. More widely, this work contributes to the growing body of work on cataloging and isolating judge-bias effects.
@@ 1 Introduction
As language models see increasing deployment as evaluators, raters, and proxies for human preference, researchers have observed consistent patterns of unreliability (Gu et al., 2025; Li et al., 2024). These patterns are frequently categorized as biases (Dietz et al., 2025; Ye et al., 2024) and are often interpreted through anthropomorphic tropes or linked to emerging theory-of-mind frameworks.
One notable example comes with the recent fascination with self-preference bias, or narcissism, of LLM evaluators (Panickssery et al., 2024; Liu et al., 2024b). A suite of recent experimental findings claim that when language models are used to compare responses to a query, the distribution of votes will skew towards outputs generated by that same model. If true, this behavior would pose critical implications for safety and alignment research; this form of situational awareness (Laine et al., 2023) risks contaminating post-training, guardrail construction, and monitoring efforts that bootstrap human preferences using LLM evaluators.
However, self-preference signals remain confounded by plausible baselines. For instance, early tests did not compare self-preference rates against ground truth accuracy, potentially mistaking a model’s accurate assessment of its own superior performance for illegitimate bias. Recent efforts now control for this by using oracle labels (Chen et al., 2025a; Wataoka et al., 2025), but still fail to disentangle the relationship between task accuracy and evaluation accuracy; that is, the ability of a judge model to assign accurate preference ratings regardless of its own skill in a given domain. The capabilities a model exhibits on a given task may inform its preferences on that same task; specifically, one might imagine that an LLM-judge’s distribution of preferences will differ depending on if it could succeed on the query in question. If evaluations on these “harder” queries show greater uncertainty, then a model might choose its own, incorrect response at random. The inability to distinguish between a model’s general evaluative proficiency and an active self-bias undermines the causality of previously established results.
@@ 2.1 Self-Preference and Other Biases in LLM Evaluators
Literature on LLM evaluation has rapidly expanded in recent years, with numerous surveys cataloging the diverse methodologies and challenges (Zheng et al., 2023; Shen et al., 2023; Chen et al., 2024; Li et al., 2024, 2025a; Gu et al., 2025). A significant subset of this work focuses on the reliability and biases of LLM evaluators, identifying various failure modes including inconsistency, susceptibility to manipulation, and context sensitivity (Koo et al., 2024; Ye et al., 2024; Dietz et al., 2025).
@@ 3.3 The Need for an Evaluator Quality Baseline
If the ILSP term controls bias, how can we characterize the distribution that produces it? All examples which fall into this subset are examples where a judge model produced a verifiably incorrect or relatively inferior response. This subset introduces a covariate on evaluations based on the judge’s task-related capability. We assert that this deficit in capability influences the voting behavior of the judge. This is analogous to a student’s performance on a multiple choice exam (pairwise preference), which may be lower-bounded by their performance in a free response setting (generative task). This student’s “multiple-choice” behavior may manifest as pure aleatory uncertainty or deference to shallow heuristics. Research on uncertainty quantification has shown that while evaluator errors do depend on task accuracy (Kalai et al., 2025; Du et al., 2024; Sychev et al., 2025), the mechanisms which drive such errors vary by model, task, and context.
@@ 4.2 Reproducing Previous Experiments
4.2.1 Verifiable Outputs (Chen et al., 2025a)
We reproduce the verifiable reasoning experiments of Chen et al. (2025a), which isolate self-preference on tasks with objectively correct answers. Datasets include MATH-500 (Hendrycks et al., 2021b), MBPP-Plus (Austin et al., 2021), and MMLU (Hendrycks et al., 2021a). Ground truth is established via execution-based evaluation for code (pass@1) and exact string matching for mathematics and multiple-choice responses. This setup enables clean separation of LSP from ILSP examples.
@@
4.2.2 Debiased Gold Judges (Chen et al., 2025b)
We reproduce the DBG-scoring methodology of Chen et al. (2025b), which evaluates self-preference on open-ended tasks including helpfulness with AlpacaEval (Li et al., 2023), truthfulness with TruthfulQA (Lin et al., 2022), and translation quality with the German-English WMT19 dataset (Barrault et al., 2019).
@@
4.2.3 Human Annotations (Panickssery et al., 2024)
We reproduce the summarization experiments of Panickssery et al. (2024), which established a correlation between self-recognition and self-preference on news summarization datasets, namely CNN/DailyMail (See et al., 2017) and XSUM (Narayan et al., 2018). Similar to DBG-Score, ground-truth summaries are generated at-scale with a “debiased gold judge”—in this case, GPT 5.2—with a prompt first tuned by small-scale human validation before tested on a larger, holdout set. We take log-probabilities through the OpenAI API, following the hyperparameters used initially. Our proxy selection matches the outcome-based approach used for other experimental setups, except that we only take a subset of potential proxies due to costs of accruing gold-judge labels from proprietary models.
@@ 4.2.4 Authorship Obfuscation
Lastly, we replicate the setup of Mahbub and Feng (2025), which investigates self-preference under stylistic perturbation using answer extraction as the verifiability signal. This approach tests whether synonym replacement and paraphrasing eliminate self-preference by disrupting surface-level cues. This work tests self-preference on the multiple-choice reading comprehension dataset QualiTY (Pang et al., 2022). Since reading comprehension responses are in multiple-choice format, we can gather ground truth labels by parsing the corresponding answer and outcome-match queries accordingly.
@@ 5.1 The Evaluator Quality Baseline Substantially Reduces Measured Self-Preference
Our baseline complements rather than contradicts prior findings: self-preference exists, but the prioritization of at-risk models has been misplaced by measurement error. Rankwise model bias orderings shift considerably after applying the Evaluator Quality Baseline. Using Alpaca Eval as an example, Llama 3.1-8B showed the highest initial bias rate of 49% before the baseline, and now shows just 3.2%—the second lowest. On the other hand, for code generation, Qwen2.5-72B went from the third-to-least to the highest rankwise bias relative to its baseline. For data-based mitigation strategies like steering vectors (Roytburg et al., 2025; Ackerman and Panickssery, 2024), having controlled measurements determines choice of mechanism design and training targets, enabling more effective interventions.
@@ Appendix A Experimental Configuration Tables
Since we are using a pairwise setting, to account for positional bias [Pezeshkpour and Hruschka, 2023] we prompt the judge twice, once where the judges prompt is first and once where the judges prompt appears second. We collect from both prompts logprobs and average them to get the probabilities for self.